Find the general solution of the equation:


π‘₯^2𝑦" βˆ’ 9π‘₯𝑦′ + 25𝑦 = 0

1
Expert's answer
2022-03-20T13:20:31-0400

"x=e^u"

"a_0\lambda(\lambda-1)(\lambda-2)...(\lambda-n+2)...+a_{n-2}\lambda(\lambda-1)+a_{n-1}\lambda+a_n=0"

"(\lambda-1)\lambda-9\lambda+25=0"

"\lambda^2-10\lambda+25=0"

"\lambda_{1,2}=5"

k=2

"y=(C_1u+C)e^{5u}"

"y=C_1x^5ln(x)+Cx^5"